OpenStudy (pumpkin1102):

8(2x – 14) + 13 = 4x – 27 What value of x makes the equation true?

OpenStudy (jdoe0001):

distribute 8, then solve for "x"

OpenStudy (warfare432):

Solution 1 Expand 16x−112+13=4x−27 2 Simplify 16x−112+13 to 16x−99 16x−99=4x−27 3 Add 99 to both sides 16x=4x−27+99 4 Simplify 4x−27+99 to 4x+72 16x=4x+72 5 Subtract 4x from both sides 16x−4x=72 6 Simplify 16x−4x= to 12x 12x=72 7 Divide both sides by 12 x=72 / 12 8 Simplify 72 / 12 to 6 x=6

OpenStudy (warfare432):

Answer = x=6

OpenStudy (warfare432):

Please mark as best response